I tried this by myself, but it didn't work out. Here is the question and the way I set it up.
If Jolly Janet has 0.5% saline to mix with 7.5% saline, how much of each is needed to get 200L of 1.0% saline?
I set it up like this:
.005X + .075(200-X)=.001(200)
I know this is wrong, but could you tell me the correct way to set it up?
Thanks.

Answer by Paul(988)

YOU always put the middle number on the right side
75x-200=5(200-x)
75x-200=1000-5x
80x=1200
x=15
200-15=185

Hence, for 7.5% about 15L is needed and for 0.5% about 185L is needed.
PAUL.